Circular No. 8345

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004bw IN MCG +00-38-19 Further to IAUC 8340, M. Moore and W. Li report the LOSS discovery

of a supernova on unfiltered KAIT images taken on May 26.39 (mag ∼ 17.8) and 27.35 UT (mag ∼ 17.6). SN 2004bw is located at α = 15^{h}03^{m}46^{s}.75, δ = −3^{o}18′16′′.4 (equinox 2000.0), which is 22′′.1 east and 7′′.2 south of the nucleus of MCG +00-38-19. A KAIT image taken on May 20.38 showed nothing at this position (limiting magnitude 19.5).

R. J. Foley, D. S. Wong, M. Ganeshalingam, W. Li, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range only 400–690 nm), obtained on May 27 UT with the Shane 3-m telescope at Lick Observatory, shows that SN 2004bw is of type I, possibly Ic, but the exact classification is uncertain due to the limited wavelength coverage. The spectrum resembles that of SN 2004aw as described by Matheson et al. (IAUC 8311).

LL ANDROMEDAE E. Waagen, AAVSO, reports that P. Schmeer (Bischmisheim, Ger-

many) has found the SU UMa-type cataclysmic variable LL And to be in superoutburst, as indicated by the following CCD magnitudes: May 19.464 UT, [16.5; 20.461, [18.0; 21.462, [18.0; 22.463, 12.6. H. Yamaoka, Kyushu University, forwards the following visual magnitude estimate of LL And by T. Ohshima (Nara, Japan): May 25.75, 13.5. Waagen adds that LL And is not a well-studied star, with this being only its third known outburst, the last outburst having occurred on 1993 Dec. 7, when it reached visual mag 13.9 and was brighter than mag 15 until Dec. 17; its usual minimum magnitude is fainter than 17.0.

SUPERNOVAE 2004bu AND 2004bv Foley et al. add that inspection of CCD spectra, taken as above, show

that SN 2004bu (IAUC 8343) is of type I, possibly resembling the peculiar SN 1998bw and SN 2002ap (Foley et al. 2003, PASP 115, 1220); however, the wavelength coverage is more limited than usual (owing to a mechanical problem in the red arm of the double spectrograph) and is insufficient for a strong confirmation. SN 2004bv (IAUC 8344) is of type Ia near maximum brightness; the weak Si II 615-nm feature and prominent Fe III lines resemble those of SN 1991T (Filippenko et al. 1992, Ap.J. 384, L15). Lacking view of the Ca II H-and-K region, it is possible that the spectrum might also resemble that of SN 1999aa (Li et al. 2001, Ap.J. 546, 734).

2004 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green